An open macroeconomic model for a hypothetical economy is represented as follows
Y= C0 +Io+Go+X0-M, M=mo+m1yd,C=co+c1yd, T=tY and Yd=Y-T
a. Show that equal change in tax and government expenditure are expansionary to the economy
b. Derive the equilibrium level of savings in the economy above
c. Derive the investment multiplier

Kindly answer me above.
1
Expert's answer
2015-03-04T08:50:16-0500
a. The equal change in tax and government expenditure are expansionary to the economy, because if the government expenditure increases (Go to G1), the GDP will increase too, as Y= C0 +Io+Go+X0-M. So, the equal change in tax and government expenditure will have expansionary effect.
b. In the equilibrium savings are equal to investment, so in our case the equilibrium level of savings is S = Io.
c. Investment multiplier is simply the multiplier effect of an injection of investment into an economy.
The investment multiplier in our case will be mi = 1/(1 - c) = 1/(1 - c1).